OpenStudy (anonymous):

17. Simplify the following: 4x – 2 – 3x 4x – (2 – 3x)

OpenStudy (anonymous):

x-2 7x-2

OpenStudy (sasogeek):

4x-2-3x = x-2 4x-(2 -3x) = 4x-2+3x =7x-2

OpenStudy (anonymous):

x-2 7x-2

OpenStudy (anonymous):

Thank you!!